Find the volume generated if the function y= square root of 1-x^2 is revolved around the x-axis (use only the positive area)


1
Expert's answer
2021-05-24T04:12:46-0400

The volume of Revolution about "O_x" is given by:


"V = \int^{x=a}_{x=b}\pi y^2dx"

So for this problem, Nothing that "1-x^2 = 0 \implies x = \pm 1" and that by symmetry we can double the volume for the region but we have to consider only positive coordinate.


"V = \int_{0}^{3} \pi(\sqrt{1-x^2})^2dx"

"V = \pi \int_{0}^{3}(1-x^2)dx"

"V = \pi[x-\dfrac{x^3}{3}]_{0}^{1}"

"V = \dfrac{2\pi}{3}"